OpenStudy (anonymous):

Which expressions would complete this equation so that the equation has no solution? 6(4x – 3) – 10x = Choose exactly two answers that are correct. A. 14x – 15 B. 2(7x – 10) C. 15x – 14 D. 2(7x – 9)

OpenStudy (anonymous):

explain please @zimmah

OpenStudy (anonymous):

i'd first simplify the expression to make it more clear

OpenStudy (anonymous):

and how do i do that can u u break it down as much as possible

OpenStudy (anonymous):

6(4x – 3) – 10x = distribute the 6 first 6*4x = 24x 6 * (-3) =-18 24x-18-10x subtract the 10x from 24x 14x-18

OpenStudy (anonymous):

ok i kinda get it can u keep going

OpenStudy (anonymous):

since we know it's 14x-18 we can already see that answer A brings no solution. because 14x-18=14x-15 you can't find any x that will work. The lines are parallel to each other and will never cross

OpenStudy (anonymous):

so cancel out a

OpenStudy (anonymous):

answer C however is not correct because the lines will cross so you can find exactly 1 solution to it. so between B and D there's another correct answer

OpenStudy (anonymous):

no, A is a correct answer, remember the questions is asking you to find exactly those ones that do not have a solution

OpenStudy (anonymous):

but it says only two answers

OpenStudy (anonymous):

yes, A is an answer and either B or D is one

OpenStudy (anonymous):

can u help with one more i suck at math please @zimmah

OpenStudy (anonymous):

2(7x – 10) = 14x-20 these also never touch just like A

OpenStudy (anonymous):

so A and B both have no solutions C has one solution and D has infinite solutions

OpenStudy (anonymous):

so its c and d